# Estimation of motion blur kernel parameters using regression convolutional neural networks

### v2

@@ 5.3.3 Error Ratio Comparisons
We calculated SSD error ratios for our proposed blur kernel estimation as well as those from Levin et al. [11] and Krishnan et al. [5] as seen in Figure 7. Recall that an error ratio of one indicates that deblurring with the estimated kernel results in an identical image to that deblurred with the ground truth kernel. An error ratio of two or higher is considered unacceptable as it was shown in [24] that such SSD error ratios indicate the presence of significant perceptually noticeable distortions in the latent image estimated using the estimated kernel. An error ratio less than 1 indicates that the image deblurred with the estimated kernel achieves a better match to the true sharp image than deblurring with the true kernel.
(a) SSD error ratio for the A3 dataset
(b) SSD error ratio for the L5 dataset
Figure 7: Error ratios for the A3 sub-dataset and L5 sub-dataset. Note that the bin at 1.0 includes error ratios $\leq 1.0$ and the bin at 3.0 includes error ratios $\geq 3.0$.
@@
We note that we have the highest cumulative error ratio at one for both A3 and L5 datasets and that our method yields an error ratio of 1.25 or less for all images in the A3 and L5 test sets. This means that our model is able to predict more accurate kernels compared to the other methods. Since our method creates the kernel using linear parameters we are less prone to additive noise in the kernel. Levin et al.’s [11] kernel prediction has noise added in the kernel since many of the pixels that are supposed to be zero are instead small numbers close to zero. This noise can be seen to affect its results in kernel prediction. Krishnan et al. [19] thresholds the small elements of the kernel to zero which increases robustness to noise.
@@ 6 Conclusions
In this paper, we have studied in detail the limitation in representation of linear blur kernels, particularly for smaller length blurs. We noticed an entanglement between length and angle when it comes to representation of the two parameters, meaning that developing a dataset that is balanced in both length and angle is difficult. Much of the research done in linear blur has implicitly assumed square, odd-sized kernels. Exploring different sized kernels gives a complete implementation of possible linear motion blur kernels that can be discovered in natural blurred images.
The complete exploration of linear motion blur gives us an advantage of training a regressive deep learning model instead of a classification model as other deep learning methods implement. With regression we can estimate more accurate blur kernel parameters which are not limited by the model needing a prior predictive kernel size. Our robustness to noise has exceeded that of other methods, even testing to a higher level than most other methods. In particular, we see about a $10\%$ drop in the $R^{2}$ metric for a $10\%$ Gaussian noise. We note, however, that the model trained on 10% Gaussian noise was robust to noise levels less than 10%, indicating that single model can serve across a wide range of noise scenarios.
In future work, we will use this exploration of linear motion blur kernels as a baseline and foundation for spatially-varying blurs. We will expand this to a patch level to decompose a spatially-varying blur image into a superposition of locally uniform blur patches. This would give us a locally linear blur field to better model atmospheric turbulence in imagery.

### v3

@@ 5.3.3 Error Ratio Comparisons
We calculated SSD error ratios for our proposed blur kernel estimation as well as those from Levin et al. [14], Krishnan et al. [9], and Carbajal et al. [25] as seen in Fig. 7. Recall that an error ratio of 1 indicates that deblurring with the estimated kernel results in an identical image to that deblurred with the ground truth kernel. An error ratio of 2 or higher is considered unacceptable as it was shown in [7] that such SSD error ratios indicate the presence of significant perceptually noticeable distortions in the latent image estimated using the estimated kernel. An error ratio $<1$ indicates that the image deblurred with the estimated kernel achieves a better match to the true sharp image than deblurring with the true kernel.
Figure 7: Error ratios for the (a) A3 sub-dataset and (b) L5 sub-dataset. The bin at 1.0 includes error ratios $\leq 1.0$ and the bin at 3.0 includes error ratios $\geq 3.0$. The results are presented for the proposed method ( ), Krishnan et al. [9] ( ), Levin et al. [14] ( ), and Carbajal et al. [25] ( ).
@@
Our proposed method has the highest cumulative histogram values for the error ratio at 1 for both A3 and L5 datasets and yields an error ratio of 1.25 or less for most images in the A3 and L5 test sets. This means that our model is able to predict more accurate kernels compared to the other methods. Since our method creates the kernel using linear parameters, we are less prone to additive noise in the kernel. Levin et al.’s [14] kernel prediction has noise added in the kernel since many of the pixels that are supposed to be zero are instead small numbers close to zero. This noise can be seen to affect its results in kernel prediction. Krishnan et al. [9] thresholds the small elements of the kernel to zero which increases robustness to noise. Carbajal et al. [25] has competitive performance for error ratios $\geq$1.25; the method is at a disadvantage at longer blur lengths due to its limitation to $33\times 3$ 33 kernels as noted above which may contribute to the diminished performance at the lowest error ratio.
Figure 8: Deblurred images for image blurred with a kernel of length 5 and angle 17. All images are deblurred with the EPLL method [30]. (a) Image deblurred using kernel estimated with our proposed method, error ratio 1.2858. (b) Image deblurred using kernel estimated with the method in Krishan et al. [9], error ratio 1.1320. (c) Image deblurred using kernel estimated with the method in Levin et al. [14], error ratio 1.5607. (d) Image deblurred using kernel estimated with the method in Carbajal et al. [25], error ratio 1.0396. (e) Original sharp image. (f) Image deblurred with the ground truth kernel. (g) Blurred image.
Figure 9: Deblurred images for sharp input image. All images are deblurred with the EPLL method [30]. (a) Image deblurred using kernel estimated with our proposed method, error ratio 1.5813. (b) Image deblurred using kernel estimated with the method in Krishan et al. [9], error ratio 1.7345. (c) Image deblurred using kernel estimated with the method in Levin et al. [14], error ratio 4.4847. (d) Image deblurred using kernel estimated with the method in Carbajal et al. [25], error ratio 1.3388. (e) Original sharp image. (f) Image deblurred with the ground truth kernel.
Figures 8 and 9 show two qualitative examples of images deblurred using the kernels estimated with our proposed method and the methods in [9, 14, 25]. The example in Fig. 8 uses a blurred image as an input and the example in Fig. 9 has a sharp image as an input for demonstration of the methods in the absence of blur. We note similar qualitative results between the deblurred images using our estimated blur kernel and the blur kernel estimated by the method in [9, 25], and that both methods yield similar results to the image deblurred with the truth kernel. We do note, however, some ringing in the images deblurred using the kernel from [9]; this is particularly noticeable in the areas around the upper power line in the image in Fig. 8(b) and the darker snow shadows in Fig. 9(b). The blur kernel estimated by the method in [14] introduces significant artifacts, especially apparent in Fig. 9(c) but also apparent in Fig. 8(c) as dark regions in the sign. The blur kernel estimated by [25] yields very good qualitative results, perhaps indicating the advantage of a kernel that is not constrained to linear motion blur when considering very large image regions. Overall, the proposed method of kernel estimation appears to yield deblurred images close to that which could be achieved with the ground truth kernel which validates this approach as the foundation for future work in estimation of motion-blur parameters in spatially-varying blur.
@@ 6 Conclusions
In this paper, we have studied in detail the limitation in representation of linear blur kernels, particularly for shorter blur lengths. We find an interdependence between length and angle in representing blur kernels, meaning that developing a dataset that is balanced in both length and angle is not possible. Furthermore, while much existing research in linear blur prediction has implicitly assumed square odd-sized kernels, we relax this assumption and allow for non-square even-sized kernels. A blurred dataset was created from the 2014 COCO dataset using a suite of blur kernels for length $r\in[1,100]$ and angle $\phi\in(-90,90]$, providing a more comprehensive variety of blur kernels than previously studied.
This thorough definition of linear motion blur and the development of a blurred dataset allowed us to train a regressive deep learning model instead of a classification model as other deep learning methods implement. With regression we can estimate more accurate blur kernel parameters which are not limited by the model needing a priori information about the kernel size. We demonstrate excellent performance in estimation of blur kernel parameters with a coefficient of determination $R^{2}\geq 0.89$ for both length and angle prediction. The robustness of the estimation to additive noise was studied for a wider range of noise than previously considered, $\sigma^{2}\in\{0.001,0.01,0.1\}$, corresponding to SNR values of $\{30,20,10\}$ dB. The regression CNN was found to be very robust to noise, with a $10\%$ drop in the $R^{2}$ metric for a $10\%$ Gaussian noise (which we note is an order of magnitude larger than the noise level of 1% commonly studied). We further note that the model trained on 10% Gaussian noise was robust to noise levels less than 10%, indicating that single model can serve across a wide range of noise scenarios. Using the estimated blur kernels in a non-blind deblurring method, we find sum of SSD error ratios of 1.25 or less for most test images, significantly outperforming the MAP-based comparison methods and competitive with a recent deep-learning based method.
In future work, we will use this exploration of linear motion blur kernels as a baseline and foundation for spatially-varying blurs. We will expand this to a patch level approach to decompose a spatially-varying blur image into a superposition of locally uniform blur patches. Extending the approach to non-uniform blur will provide a better means to estimate the spatially varying blur resulting from atmospheric turbulence in images. Additionally, it would be interesting to directly compare our approach to other deep learning approaches which either directly estimate length and angle or which assume locally linear motion blur kernels, e.g., [22, 24, 15, 26].
Disclosures
The authors have no conflicts of interest to disclose.